# PULASki: Learning inter-rater variability using statistical distances to improve probabilistic segmentation

### v1

@@ 4.3 Training and Evaluation
For evaluating the results, Generalised Energy Distance (GED) szekely2013energy ; kohl2018 was employed, calculated using the following equation:
@@
Although GED serves as a metric to evaluate how close the distribution of the prediction is to the distribution of the annotations, it does not tell anything regarding whether or not the variability present in the annotation is similar to the variability in the predictions produced by the method. Krippendorff’s alpha ($K\alpha$) krippendorff2018content is a statistical measure that shows the agreement among the samples - commonly used to evaluate the inter-rater agreement. Hence, $K\alpha$ was used as an additional metric to evaluate how closely the methods reproduced the inter-rater disagreement, and is calculated as:
@@
Finally, the statistical significance of the variance noted across various methods and metrics was ascertained utilising the signed Wilcoxon Rank test wilcoxon1992individual .
@@ 5.1 Statistical measures of predicted segmentation variability
PULASKi with the Hausdorff divergence shows the best performance across both experiments, in terms of both GED and $K\alpha$. It provides an improvement over the standard Probabilistic U-Net of 56 and 51 % for 2D and 3D vessel segmentation and 46 and 33 % for 2D and 3D MS lesion segmentation, in terms of the average GED. The chosen statistical distances with PULASKi more often that not provide a statistically significant improvement in terms of both evaluation metrics over baselines at the 5% significance level, particularly in the vessel segmentation task (see Table 2). PULASKi (Hausdorff) was the only method to consistently outperform baselines at the 5% significance level. Amongst the baselines methods, MCDO with FTL loss allows for the best representation of inter-rater variability in terms of both GED (average of 0.45 for 2D and 0.49 for 3D in vessel segmentation and 0.47 for 2D and 0.55 for 3D in MS segmentation) and $K\alpha$, rather surprisingly. VI-MH U-Net provides a slight improvement over the Probabilistic U-Net in the 2D experiments. The SSN was completely unable to represent the underlying variability in the training data, as evidenced by the large GED and Krippendorff alpha scores. We posit that this is due to the cross-entropy term style loss function being bad for class imbalanced data and due to the lack of regularisation, which is present in PULASKi, Probabilistic U-Net and VI-MH U-Net.
@@
Table 2: $p$-values obtained using the Wilcoxon rank test while comparing the different PULASki methods against the baselines and annotations combining both data sets. Significance indicated at $\alpha=0.05$ (✓: Significant, ✗: Not Significant).
Prob UNet MCDO UNet SSN VIMH Annotation
@@
GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$
Hausdorff 2D ✓($<10^{-6}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<0.001$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<10^{-7}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<10^{-6}$) ✓($<10^{-4}$) ✓($<10^{-4}$) $(0.13)$ $(1.0)$
Sinkhorn 2D ✓($<0.001$) ✓($<10^{-4}$) ✓($<10^{-4}$) $(1.0)$($<10^{-4}$) ✓($<10^{-4}$) ✓($<10^{-7}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<0.01$) ✓($<10^{-4}$) ✓($<0.001$) ✓($<10^{-4}$) ✓($<10^{-4}$)
Hausdorff 3D ✓($<10^{-5}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<0.01$) ✓($<10^{-4}$) ✓($<10^{-4}$) - - - - - - ✓($<10^{-4}$) ✓($<10^{-4}$)
@@ 5.2 Qualitative measures of predicted segmentation variability
While evaluation metrics provide a concise quantitative measure of how closely the predicted distribution matches the annotations, qualitative results can be useful to determine which specific features are well represented by different methods. We plotted the frequency of labelling as vessel or lesion at a voxel level across 10 predicted segmentations from each method (Rate of Occurrence (RoO) see Figure 5. The brighter (yellow), the voxel appears, the more often it is labelled as a vessel or MS lesion throughout the output plausible segmentations. In line with the quantitative metrics, the PULASKi approach generates segmentations with structure and levels of agreement between annotations that are more consistent with the training data (see Figure 3), for both vessel and lesion segmentation. However, frequencies appear to be better represented in the vessel segmentation task than in the MS lesion task, where the extent of lesions is smaller and level of agreement across predicted segmentations higher than in the training data (see Figure 3(B)). This is true not just for PULASKi, but across all methods and could be attributed to the lower inter-rater agreement in the annotated MS data. Finally, vessel thickness and lesion extents are slightly over-estimated in the PULASKi method in 2D slice based training compared to 3D volumes.
@@ 5.3 Pixel-wise Probability Distribution
Figure 9: Comparison Maximum Intensity Projection (MIP) generated from the median pixel-wise probability maps for a given subject in 2D vs 3D Hausdorff implementation for the vessel segmentation task. Results are shown for threshold of 0.5 in (A) and 0.001 in (B). The green circle encompasses a region that includes small vessels, and its sensitivity to selected thresholds is particularly pronounced in the 2D implementation.
Training models to segment inherently 3D volumes using training data consisting of 2D slices, whilst computationally cheaper, generally produces poorer quality segmentations. As described in Section 5.1, GED scores were comparable between 2D and 3D implementations of the best performing method (PULASki with Hausforff distance), while the $K\alpha$ value was slightly lower for 3D compared to 2D. However, 3D mesh visualisation of the segmented vessels (Figure 7) reveals unrealistic “stacked” or “stapled” regions more prevalent in the 2D implementation than 3D. Moreover, it can be observed in Figure 8 that the pixel-wise probability of a voxel being a vessel exhibits a nearly discrete distribution of probability in 2D, whereas in 3D it gradually decreases from the centre to the edges. This can be attributed the lack of continuity between 2D slices, which is much better preserved when training on 3D patches. In the 3D implementation, predictions during inference are generated in overlapping patches. The final prediction is obtained by averaging over the overlapping region, thereby providing a better transition across patches. This makes the 3D results better suited for downstream tasks, e.g., 3D surface generation for CFD or DA, where smooth transitions are important. This smoothness can further be observed in figure 6, where pixel-wise probabilities have a much more spatially consistent structure in 3D than 2D. However, the averaging operation that allows smoother transitions in the 3D implementation has the potential to reduce variability in the generated segmentations, as indicated in Figure 5 and the higher $K\alpha$ values (Table 1). Moreover, the 3D implementation yields more certain pixel-wise probabilities. The Maximum Intensity Projection (MIP) of the pixel-wise probabilities for all values that exceed a probability of 0.5 (top row) and 0.001 (bottom row) confirm greater certainty and stability in the 3D implementation (Figure 9. Specifically, for small vessels (green circle), the 2D projection represents the area only for the lower threshold as a connected vessel segment, whereas the 3D implementation illustrates analogous structures for both thresholds. Although smaller vessels appear adequately represented in the lower threshold of the 2D segmentation, larger vessels appear to be overestimated.
@@ 5.6 Methodological insights
The superior performance of PULASKi with an appropriate statistical distance compared to baselines can be primarily attributed to two factors; 1) effective learning of a low dimensional representation on which uncertainty quantification can be performed efficiently and 2) the chosen distribution based reconstruction losses that better capture inter grader variability (that is, to better learn $p(y|z,x)$). A significant challenge in the case studies considered here is class imbalance; for the MS training data only 0.13% of the full training data set corresponds to lesions and 2.56% corresponds to vessel in the intracranial vessel segmentation data on average. PULASKi is more robust to class imbalance issues than the adopted baselines, as no specific measures were adopted to deal with it while the original formulation of the Probabilistic U-Net failed to even converge during training until the cross-entropy was replaced with the Focal Tversky Loss (FTL) abraham2019novel .
The chosen statistical distances in PULASKi are also numerically better suited to comparing the distribution of segmentations for a given image with the model output distribution. Empirical cross-entropy style losses used in the baselines poorly capture the variability in the conditional distribution $p(y|x)$, as the log probabilities for all segmentations, regardless of the corresponding image, are lumped together through a sum (see e.g., 12). Poor performance of the SSN can be attributed to similar issues, given the similarity of the loss function to that of the Probabilistic U-net, although there are likely additional shortcomings of the Deepmedic architecture for the tasks at hand. The lower resolution input might ignore smaller structures - a very frequent occurrence in both tasks performed here. Thed We attribute improved performance of the PULASKi using the Hausdorff divergence, over the theoretically more accurate de-biased Sinkhorn divergence to potential instabilities introduced when solving for the more complicated optimal transport loss in (32). The use of multiple “heads” in the MH VI-Unet appears to not be flexible enough to represent inter-rater variability in the examined case studies, although this could potentially be improved by increasing the number of heads. This would also then increase the computational expense.
@@ 6 Discussion
Segmenting complex vessel geometries and tissue lesions from noisy MRI remains a challenging task, despite improvements in imaging technology. Often times there is significant variation in labels from human annotators for the same image. Supervised and semi-supervised deep learning based segmentation is rapidly developing to provide more accurate results in a fraction of the time of traditional methods. Ensuring these methods faithfully capture the underlying uncertainty in labelled data is paramount, so as to prevent over-confident diagnoses of diseases and model predictions. Our method provides a significant improvement on many existing methods for capturing inter-rater variability in deep learning based medical image segmentation, for the same amount of training data. It is easily implementable and can be used to provide practitioners with pixel-wise uncertainties and generate several whole volume plausible segmentations for ensemble-based modelling purposes involving computational fluid dynamics models and data assimilation gaidzik2021hemodynamic .
@@
Experimental results on intracranial image segmentation tasks show that the distribution based loss functions utilised in our proposed method, PULASKi, drastically improve the ability to learn diverse plausible segmentations for a given image. The loss functions are also far less sensitive to class imbalance in the training data set, based on experiments in both vessel and multiple sclerosis segmentation tasks in 2D and 3D cases. Furthermore, there are minimal extra hyper-parameters to be specified compared to existing methods (e.g., probabilistic U-net), whilst simultaneously providing dramatic improvement in performance. Our experimental results also demonstrate a significant improvement in both the anatomical plausibility of generated segmentations and representation of uncertainty when training on 3D patches rather than 2D slices. In particular, vessel thickness seems to be overestimated easily in the 2D implementation with an uncertain threshold and vessel boundaries are not as smooth as in 3D.
@@
The representation of small vessels is drastically improved using 3D training data, a task that has traditionally been incredibly challenging. Manual and semi-automated segmentation, despite being time-consuming, is typically deemed reliable primarily for medium to large vessels, as indicated in previous studies jerman2015 ; Chatterjee2020 . The proposed work is one of the first studies to demonstrate reliable segmentation of small vessels and supports the results of Chatterjee2020 by producing a set of plausible segmentations. Moreover, PULASki was the only method showing variability in the subset of possible segmentations within the area of smaller vessels, as illustrated in the RoO results (Figure 5) This also aligns more closely with the annotated data, as depicted in 3. This provides significant opportunity to improve the modeling of cerebral small-vessel which supports the research in diseases that are difficult to predict and manage, e.g., CSVD. In the MS segmentation task, PULASKi also provides an improved representation of inter-rater variability over the baselines, although “outlier” or “extreme” annotations are still not represented well. This suggests potential for further improvements to quantify distributional tails more accurately. 5
@@
While PULASki, as well as the baseline methods presented here, are aimed at providing multiple plausible segmentations. However, there can be situations where obtaining one single segmentation per input is required (e.g. at clinics). In such scenarios, majority voting from the predicted plausible segmentations can be taken, or even the segmentation from the prior mean can be utilised.
@@
It should be noted that the distribution-based loss functions in PULASKi are more computationally expensive to evaluate than the standard cross-entropy or FTL used in the benchmark methods. However, we anticipate that this is generally not prohibitive, as training is typically undertaken offline and there is no further computational burden for inference tasks, compared to other variational autoencoder based methods.

### v2

@@ 4.3 Training and Evaluation
For evaluating the results, Generalised Energy Distance (GED) [57, 25] was employed, calculated using the following equation:
@@
Although GED serves as a metric to evaluate how close the distribution of the prediction is to the distribution of the annotations, it does not tell anything regarding whether or not the variability present in the annotation is similar to the variability in the predictions produced by the method. Krippendorff’s alpha ($K\alpha$) [58] is a statistical measure that shows the agreement among the samples - commonly used to evaluate the inter-rater agreement. Hence, $K\alpha$ was used as an additional metric to evaluate how closely the methods reproduced the inter-rater disagreement, and is calculated as:
@@
Finally, the statistical significance of the variance noted across various methods and metrics was ascertained utilising the signed Wilcoxon Rank test [59].
@@ 5.1 Statistical measures of predicted segmentation variability
PULASKi with the Hausdorff divergence shows the best performance across both experiments, in terms of both GED and $K\alpha$. It provides an improvement over the standard Probabilistic U-Net of 56 and 51 % for 2D and 3D vessel segmentation and 46 and 33 % for 2D and 3D MS lesion segmentation, in terms of the average GED. The chosen statistical distances with PULASKi more often that not provide a statistically significant improvement in terms of both evaluation metrics over baselines at the 5% significance level, particularly in the vessel segmentation task (see Table 2). PULASKi (Hausdorff) was the only method to consistently outperform baselines at the 5% significance level. Amongst the baselines methods, MCDO with FTL loss allows for the best representation of inter-rater variability in terms of both GED (average of 0.45 for 2D and 0.49 for 3D in vessel segmentation and 0.47 for 2D and 0.55 for 3D in MS segmentation) and $K\alpha$, rather surprisingly. VI-MH U-Net provides a slight improvement over the Probabilistic U-Net in the 2D experiments. The SSN was completely unable to represent the underlying variability in the training data, as evidenced by the large GED and Krippendorff alpha scores. We posit that this is due to the cross-entropy term style loss function being bad for class imbalanced data and due to the lack of regularisation, which is present in PULASKi, Probabilistic U-Net and VI-MH U-Net. Both stages of D-Persona and CIMD with $T=100$ failed to produce consistent and reliable results for both MS and vessel segmentation tasks (quantitative results can be found in the supplementary materials, Fig. 12 and Tab. 6). Hence, they were discarded from further analyses. Consequently, these methods were excluded from further analyses. For MS segmentation using CIMD, $T$ was increased to the original $T=1000$ (requiring more than 10 hours per volume), which yielded results comparable to those from MCDO in terms of GED (0.47). The low Krippendorff alpha scores indicate a high degree of heterogeneity in the generated plausible segmentations. The time required to process each volume (more than 60 hours) using CIMD with $T=1000$ for vessel segmentation rendered it impractical, and thus, this task was not undertaken.
@@
Table 2: $p$-values obtained using the Wilcoxon rank test while comparing the different PULASki methods against the baselines and annotations combining both data sets (*CIMD: only for MS Segmentation). Significance indicated at $\alpha=0.05$ (✓: Significant, ✗: Not Significant).
Prob UNet MCDO UNet SSN VIMH CIMD* Annotation
@@
GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ GED $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$ $\mathrm{K\alpha_{all}}$ $\mathrm{K\alpha_{ROI}}$
Hausdorff 2D ✓($<10^{-6}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<0.001$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<10^{-7}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<10^{-6}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<0.05$) ✓($<10^{-6}$) ✓($<10^{-6}$) $(0.13)$ $(1.0)$
Sinkhorn 2D ✓($<0.001$) ✓($<10^{-4}$) ✓($<10^{-4}$) $(1.0)$($<10^{-4}$) ✓($<10^{-4}$) ✓($<10^{-7}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<0.01$) ✓($<10^{-4}$) ✓($<0.001$) ✗($0.8$) ✓($<10^{-6}$) ✓($<10^{-6}$) ✓($<10^{-4}$) ✓($<10^{-4}$)
Hausdorff 3D ✓($<10^{-5}$) ✓($<10^{-4}$) ✓($<10^{-4}$) ✓($<0.01$) ✓($<10^{-4}$) ✓($<10^{-4}$) - - - - - - - - - ✓($<10^{-4}$) ✓($<10^{-4}$)
@@ 5.2 Qualitative measures of predicted segmentation variability
While evaluation metrics provide a concise quantitative measure of how closely the predicted distribution matches the annotations, qualitative results can be useful to determine which specific features are well represented by different methods. We plotted the frequency of labelling as vessel or lesion at a voxel level across 10 predicted segmentations from each method (Rate of Occurrence (RoO) see Figure 5. The brighter (yellow), the voxel appears, the more often it is labelled as a vessel or MS lesion throughout the output plausible segmentations. In line with the quantitative metrics, the PULASKi approach generates segmentations with structure and levels of agreement between annotations that are more consistent with the training data (see Figure 3), for both vessel and lesion segmentation. However, frequencies appear to be better represented in the vessel segmentation task than in the MS lesion task, where the extent of lesions is smaller and level of agreement across predicted segmentations higher than in the training data (see Figure 3(B)). This is true not just for PULASKi, but across all methods and could be attributed to the lower inter-rater agreement in the annotated MS data. Finally, vessel thickness and lesion extents are slightly over-estimated in the PULASKi method in 2D slice based training compared to 3D volumes. CIMD, as a diffusion-based model, produces segmentations that exhibit the highest heterogeneity compared to all other models. However, in comparison to the annotations, it identifies regions as lesions that were not selected by the annotators, which can possibly be attributed to the issue of ”hallucination of lesions”.
@@ 5.3 Pixel-wise Probability Distribution
Training models to segment inherently 3D volumes using training data consisting of 2D slices, whilst computationally cheaper, generally produces poorer quality segmentations. As described in Section 5.1, GED scores were comparable between 2D and 3D implementations of the best performing method (PULASki with Hausforff distance), while the $K\alpha$ value was slightly lower for 3D compared to 2D. However, 3D mesh visualisation of the segmented vessels (Figure 7) reveals unrealistic “stacked” or “stapled” regions more prevalent in the 2D implementation than 3D. Moreover, it can be observed in Figure 8 that the pixel-wise probability of a voxel being a vessel exhibits a nearly discrete distribution of probability in 2D, whereas in 3D it gradually decreases from the centre to the edges. This can be attributed the lack of continuity between 2D slices, which is much better preserved when training on 3D patches. In the 3D implementation, predictions during inference are generated in overlapping patches. The final prediction is obtained by averaging over the overlapping region, thereby providing a better transition across patches. This makes the 3D results better suited for downstream tasks, e.g., 3D surface generation for CFD or DA, where smooth transitions are important. This smoothness can further be observed in figure 6, where pixel-wise probabilities have a much more spatially consistent structure in 3D than 2D. However, the averaging operation that allows smoother transitions in the 3D implementation has the potential to reduce variability in the generated segmentations, as indicated in Figure 5 and the higher $K\alpha$ values (Table 1). Moreover, the 3D implementation yields more certain pixel-wise probabilities. The Maximum Intensity Projection (MIP) of the pixel-wise probabilities for all values that exceed a probability of 0.1 (A), 0.01 (B) and 0.001 (C) confirm greater certainty and stability in the 3D implementation (Figure 11). Specifically, for small vessels (green and orange circle), the 2D projection represents the area only for the lower threshold as a connected vessel segment, whereas the 3D implementation illustrates analogous structures for both thresholds. Although smaller vessels appear adequately represented in the lower threshold of the 2D segmentation, larger vessels appear to be overestimated. This is particularly evident when considering the large vessels in the 0.001 threshold. They appear unrealistically large and overestimated while obscured by noise.
5.5 Anatomical plausibility of generated segmentations
@@ 5.6 Methodological insights
The superior performance of PULASKi with an appropriate statistical distance compared to baselines can be primarily attributed to two factors; 1) effective learning of a low dimensional representation on which uncertainty quantification can be performed efficiently and 2) the chosen distribution based reconstruction losses that better capture inter grader variability (that is, to better learn $p(y|z,x)$). A significant challenge in the case studies considered here is class imbalance; for the MS training data only 0.13% of the full training data set corresponds to lesions and 2.56% corresponds to vessel in the intracranial vessel segmentation data on average. PULASKi is more robust to class imbalance issues than the adopted baselines, as no specific measures were adopted to deal with it while the original formulation of the Probabilistic U-Net failed to even converge during training until the cross-entropy was replaced with the Focal Tversky Loss (FTL) [13].
The chosen statistical distances in PULASKi are also numerically better suited to comparing the distribution of segmentations for a given image with the model output distribution. Empirical cross-entropy style losses used in the baselines poorly capture the variability in the conditional distribution $p(y|x)$, as the log probabilities for all segmentations, regardless of the corresponding image, are lumped together through a sum (see e.g., 12). Poor performance of the SSN can be attributed to similar issues, given the similarity of the loss function to that of the Probabilistic U-net, although there are likely additional shortcomings of the Deepmedic architecture for the tasks at hand. The lower resolution input might ignore smaller structures - a very frequent occurrence in both tasks performed here. Thed We attribute improved performance of the PULASKi using the Hausdorff divergence, over the theoretically more accurate de-biased Sinkhorn divergence to potential instabilities introduced when solving for the more complicated optimal transport loss in (45). The use of multiple “heads” in the MH VI-Unet appears to not be flexible enough to represent inter-rater variability in the examined case studies, although this could potentially be improved by increasing the number of heads. This would also then increase the computational expense. The latest baseline employed in this work, CIMD (with $T=1000$), yielded a comparable GED to the MCDO U-Net, whilst exhibiting heterogeneity exceeding that of the ground truth for the 2D MS segmentation task. However, the inference time of this model, being significantly greater than that of all other methods, coupled with the issue of ”hallucination of lesions” identified through closer qualitative analysis, rendered the model neither particularly useful nor reliable in comparison to the others. Owing to the exceedingly high inference time, CIMD (with $T=1000$) was not utilised for 2D vessel segmentation or either of the 3D segmentation tasks.
@@ 6 Discussion
Segmenting complex vessel geometries and tissue lesions from noisy MRI remains a challenging task, despite improvements in imaging technology. Often times there is significant variation in labels from human annotators for the same image. Supervised and semi-supervised deep learning based segmentation is rapidly developing to provide more accurate results in a fraction of the time of traditional methods. Ensuring these methods faithfully capture the underlying uncertainty in labelled data is paramount, so as to prevent over-confident diagnoses of diseases and model predictions. Our method provides a significant improvement on many existing methods for capturing inter-rater variability in deep learning based medical image segmentation, for the same amount of training data. It is easily implementable and can be used to provide practitioners with pixel-wise uncertainties and generate several whole volume plausible segmentations for Monte Carlo-based modelling purposes involving computational fluid dynamics models and data assimilation [8].
@@
Experimental results on intracranial image segmentation tasks show that the distribution based loss functions utilised in our proposed method, PULASKi, drastically improve the ability to learn diverse plausible segmentations for a given image. The loss functions are also far less sensitive to class imbalance in the training data set, based on experiments in both vessel and multiple sclerosis segmentation tasks in 2D and 3D cases. Furthermore, there are minimal extra hyper-parameters to be specified compared to existing methods (e.g., probabilistic U-net), whilst simultaneously providing dramatic improvement in performance. Our experimental results also demonstrate a significant improvement in both the anatomical plausibility of generated segmentations and representation of uncertainty when training on 3D patches rather than 2D slices. In particular, vessel thickness seems to be overestimated easily in the 2D implementation with an uncertain threshold and vessel boundaries are not as smooth as in 3D. In terms of practical applications, presumably 3D implementations are generally more beneficial, particularly when smooth, continuous surfaces are required for advanced downstream tasks, such as in CFD. While 2D approaches can still be useful for gaining an initial understanding of the distribution of possible segmentations, they are not ideal for tasks that demand high accuracy and reliability. In the medical field, 2D segmentation may be sufficient where the focus is on identifying potential regions - providing a general idea - where lesions (e.g. MS or tumour) could occur, rather than needing to delineate smooth, continuous boundaries.
Figure 11: Comparison Maximum Intensity Projection (MIP) generated from the median pixel-wise probability maps for a given subject in 2D vs 3D Hausdorff implementation for the vessel segmentation task. Results are shown for threshold of 0.1 in (A), 0.01 in (B) and 0.001 in (C). The green and orange circle encompasses a region that includes small vessels, and its sensitivity to selected thresholds is particularly pronounced in the 2D implementation.
It is to be noted that training in 3D significantly increases the computational requirements. Inference using 3D patches (like done here) might result in some reduction in variability due to the averaging operation performed for the overlaps (as discussed earlier). This can possibly be avoided by performing the inference on the whole 3D volume. Given the computational limitations, we could not perform it during this research.
@@
The representation of small vessels is drastically improved using 3D training data, a task that has traditionally been incredibly challenging. Manual and semi-automated segmentation, despite being time-consuming, is typically deemed reliable primarily for medium to large vessels, as indicated in previous studies [60, 12]. The proposed work is one of the first studies to demonstrate reliable segmentation of small vessels and supports the results of [12] by producing a set of plausible segmentations. Moreover, PULASki was the only method showing variability in the subset of possible segmentations within the area of smaller vessels, as illustrated in the RoO results (Figure 5) This also aligns more closely with the annotated data, as depicted in 3. This provides significant opportunity to improve the modelling of cerebral small-vessel which supports the research in diseases that are difficult to predict and manage, e.g., CSVD. In the MS segmentation task, PULASKi also provides an improved representation of inter-rater variability over the baselines, although “outlier” or “extreme” annotations are still not represented well. This suggests potential for further improvements to quantify distributional tails more accurately.
@@
Although the focus of our work is in providing multiple plausible segmentations, there can be situations where obtaining one single segmentation per input is required (e.g. at clinics). In such scenarios, majority voting from the predicted plausible segmentations can be taken, or even the segmentation from the prior mean can be utilised.
@@
It should be noted that the distribution-based loss functions in PULASKi are more computationally expensive to evaluate than the standard cross-entropy or FTL used in the benchmark methods. However, we anticipate that this is generally not prohibitive, as training is typically undertaken offline and there is no further computational burden for inference tasks, compared to other variational autoencoder based methods. It is also to be noted that this research did not use FID loss for the 3D experiments as the loss model was 2D and this can be extended in the future by using a 3D loss calculation model trained on a 3D dataset.
Further challenges with PULASki include its ability to adequately capture the full range of distribution for both vessel and multiple sclerosis segmentation. Whilst the 3D segmentation is more robust, it nevertheless fails to encompass the entire distribution range and may omit certain outliers. However, it demonstrates closer alignment with the annotated range compared to the baseline models we evaluated.
Moreover, the current implementation can erroneously identify regions devoid of vessels as containing vessels. To mitigate this issue, we propose incorporating a consistency criterion to evaluate physical plausibility in future work. This approach would be contingent upon the specific task for which PULASki is employed; for instance, in vessel detection, the code could assess connectivity to ensure accuracy. These observations underscore areas where our implementation could be refined to improve its efficacy in segmentation tasks.